OpenStudy (bekkah323):

The body mass indices of American men between ages 30 and 50 are normally distributed with a mean of 26.2 and a standard deviation of 4.7. Several health organizations have declared that men with a BMI of 25 or greater are overweight and that men with a BMI of 30 or greater are obese. According to these values, what percentage of American men are obese?

OpenStudy (agent0smith):

Do you know how to find z-scores? Find the z-score for 30,using the given mean and standard deviation.

OpenStudy (loser66):

P(x> 30) =? just put it in calculator.

OpenStudy (reemj):

Z-score = ( value - mean ) / standard deviation P ( x > 30 ) = 1 - 0.8106 = 0.1894

OpenStudy (bekkah323):

x-m/standard deviation

OpenStudy (bekkah323):

30-26.2/4.7

OpenStudy (bekkah323):

where did you get 0.8106

OpenStudy (reemj):

I took it from z-score table http://math.tutorvista.com/statistics/z-score-table.html

OpenStudy (bekkah323):

ok I got a different answer. I took 30-26.2/4.7=.810 On the z score table that gave me 0.7910 1-0.7910= 20.9%

OpenStudy (agent0smith):

That looks right^ and reemj didn't show much work so...

OpenStudy (loser66):

yup, I got the same answer on my calculator.